I am trying to figure out how to solve polynomial equations and my instructor is not very good at explaining things so that I understand them. Can someone explain to me how to solve polynomial equations. Example:
(4x+9)(15x^2-7x-2)=0
I don't understand algebra at all, and I'm lacking the help from my instructor.

Found 2 solutions by ewatrrr, Alan3354:
Answer by ewatrrr(24785)


Hi
Solving
(4x+9)(15x^2-7x-2) =0 |the concept is finding products = 0 to solve for x
factoring further:
factoring the (15x^2-7x-2)portion: factoring: expressing as 2 products..
In this case: (_x 1)(_x 2)|the hint here is the -7x middle term
(4x+9)(5x + 1)(3x - 2)= 0 |strategically placing the 5 and 3 & ± works
Note:SUM of the inner product(3x) and the outer product(-10x) = -7x
(4x+9)(5x+1)(3x -2)= 0 |NOW we are prepared to solve for x(s)
(4x+9)= 0 x = -9/4
(5x+1)= 0 x = -1/5
(3x -2)= 0 x = 2/3
Solution set is x = {-9/4,-1/5,2/3}

Answer by Alan3354(69443)

(4x+9)(15x^2-7x-2)=0
--------------
If the product is zero, then one of the factors must be zero. See that?
If (4x+9) = 0
4x = -9
x = -9/4
--------
If (15x^2-7x-2)=0
This is not as obvious, but you can factor it:
(5x+1)*(3x-2) = 0
Then either (5x+1) = 0 --> x = -1/5
----------
or (3x-2) = 0 --> x = 2/3
--------------
The 3 solutions are
x = -9/4
x = -1/5
x = 2/3